OpenStudy (anonymous):

$299 DVD player; 7% sales tax

Directrix (directrix):

Total cost is $299 times 1.07 = $ 319.93

Directrix (directrix):

Question?

OpenStudy (anonymous):

can you do another one for me

Directrix (directrix):

Post it here.

OpenStudy (anonymous):

1. $49 tie; 15% discount 2. $59 power tool; 5% sales tax lol my bad i ment 2 question :)

Directrix (directrix):

Cost of $49 tie at 15% discount --------------- $49 times .85 = $41.65 Cost of $59 power tool with 5% sales tax -------------- 59 times 1.05 = $61.95

OpenStudy (anonymous):

doesnt 5% equal $2.50

Directrix (directrix):

Cost of $59 power tool with 5% sales tax -------------- 59 times 1.05 = 59 (1 + .05) = 59 + 2.95 = %61.95 I'm doing one-step calculations. The tool is 59 so I want all of it (100% = 1) plus 5% (5% - .05) more. A multiplier of 1.05 will do and yield the correct answer. --- You did just the discount [doesnt 5% equal $2.50] but that is not the question.

OpenStudy (anonymous):

oh okaee i get it now i have a few more feel to do them or naw

Directrix (directrix):

Post in a new thread.